A financial analyst is attempting to assess the future dividend policy of Environmental Systems by examining its life cycle. She anticipates no payout of earnings in the form of cash dividends during the development stage (I). During the growth stage (II), she anticipates 19 percent of earnings will be distributed as dividends. As the firm progresses to the expansion stage (III), the payout ratio will go up to 32 percent and eventually reach 56 percent during the maturity stage (IV).


a. Assuming earnings per share will be as follows during each of the four stages, indicate the cash dividend per share (if any) during each stage. (Leave no cells blank - be certain to enter "0" wherever required. Do not round intermediate calculations and round your answers to 2 decimal places.)

Stage I $ .35
Stage II 1.70
Stage III 2.60
Stage IV 3.50

Dividends:

Stage I:

Stage ll:

Stage lll:

Stage lV:


b. Assume in Stage IV that an investor owns 330 shares and is in a 15 percent tax bracket. What will be the investor’s aftertax income from the cash dividend? (Do not round intermediate calculations and round your answer to 2 decimal places.)


Aftertax income_____

c. In what two stages is the firm most likely to utilize stock dividends or stock splits? (Select two answers. Single click the box with the question mark to produce a check mark for a correct answer and double click the box with the question mark to empty the box for a wrong answer.)

Stage I
Stage II
Stage III
Stage IV
Answer #1

Answer a:

Cash dividend per share = Earnings per share * Payout ratio

Dividends:

Stage I = $0.3 * 0% = $0

Stage II = $1.70 * 19% = $0.32

Stage III = $2.60 * 32% = $0.83

Stage IV = $3.50 * 56% = $1.96

Answer b:

Assume in Stage IV that an investor owns 330 shares and is in a 15 percent tax bracket

Stage IV dividend per share = $1.96

Total dividends = Number of shares * dividend per share = 330 * 1.96 = $646.80

After-tax income =Total dividends * (1 - tax rate) = 646.80 * (1 -15%) = $549.78

After-tax income = $549.78

Answer c:

The two stages is the firm most likely to utilize stock dividends or stock splits are growth stage (II) and expansion stage (III).

Since in these two stages company would like to hold cash to invest in growth projects and potential investments requirements.